a sample of hydrogen gas(2.016g/mol)was found to have a pressure of 634mmhg,a temperature of 32°c and a volume of 9.15 L.how many litres of this sample of Ar occupy at a pressure of 455mmhg and a temperature of -53°c?
1
Expert's answer
2020-06-05T05:14:33-0400

pV=nRT.

n=pV/RT= 634 x 9.15 / 305 x 62400= 0.0003 .

V=nRT/p= 0.0003 x 62400 x 220 / 455 = 9,19648352.